OpenStudy (anonymous):

Solve the following inequality and write your solution using set builder notation. -3/8x<4/15

OpenStudy (anonymous):

will you explain how you got that answer please

OpenStudy (saifoo.khan):

did a mistake, \[x> - \frac{32}{45}\]

OpenStudy (saifoo.khan):

\[\frac38x > -\frac{4}{15}\] \[3x > -\frac{32}{15}\] \[x > - \frac{32}{45}\]